Back to the NIKA2 main page

Daily Reports (2026/02/10-17, NIKA2 run 79, Cryo run 91)

Observers on-site: Miren Muñoz (Tue.-Fri.), Maite Beltrán, David Taylor, Philippe André, Angel Bongiovanni (Fri. to Tue.)

AoD: Angel Bongiovanni

Operators: Gloria Liñán (night), Pedro Martínez (day)

Telescope crew and observers arrived safe. Thanks to German for his support. Lot of snow around, but melting down. The telescope looks quite clean. Raining. Ambient temp. around 0C. No wind.

Organization of shifts:
23:00-06:00 LT: Maite
06:00-13:00 LT: Philippe
13:00-20:00 LT: David
19:00-23:00 LT: Miren/Angel

Important

Focus offsets shall be measured on Ar1
Pointing offsets shall be measured on Ar2

Pool-Account usage: Accounting of observing times

Focus and Beam maps accounting

Please fill the Table with the Ar1,2,3 focus values as follows. At best, make one table per day.

Similarly, if any beam maps (total power or polarimetry) are taken:

Tuesday, 10th February 2026

@9:30-15:30 LT: Maintenance time.

@15:30-20:00: All together. Heavy cloudy and raining. Mild wind. A primer about NIKA2 observations.

No observations - adverse weather. Forecast still being unfavourable.

Wednesday, 11th February 2026

00:00-06:00: Maite. Rain and wind all night.

06:00-13:00: Philippe. It's still raining mid-morning. A possible pause in the afternoon is provided, but the rain would come back at night:

forecast_Wed

Mild <wind>: ~10 m/s. Temperature outside is slightly above 0C.

13:00-19:00: David. Rain and wind.

19:00-23:00: Miren. Frozen mix, dense fog.

23:00-24:00: Maite. Fog and rain. No observations all day.

Thursday, 12th February 2026

00:00-06:00: Maite. Heavy rain, which stopped ca. 03h; then strong wind and frozen rain. Deicing ON at 06h. No observations.

06:00-13:00: Philippe. Wind speed is dropping a bit. Sky is clearing just before dawn. Deicing OFF (08:30). Pedro takes actions for checking taumeter box, telescope surface, chiller fans, etc. Auxiliary anemometer is installed.

Dawn's colours

13:00-19:00: David. Taking over by doing a calib_1scan on NGC7027 (s59) and CRL2688 (s61). Opacity/instabilities are rising. Waiting for a while to see if conditions improve or not. At ~14h (LT), we confirm that opacity and its oscillations get worse, although the forecasts indicate the contrary.

19:00-23:00: Miren. Taking over when the last scans on Cl0024+1652 were finishing.

23:00-24:00: Maite. Low clouds and very high opacity (~1.6). Have tried different pointing on Uranus without succeeding. Shifting to project test for a while. Impossible to point on Uranus at 30 degrees elevation. We are completely inside the cloud and the wind speed is high (we have got tracking problems).

* At 24:00, the telescope is parked. The forecast indicates that the clouds are going to thicken and the wind speed to increase. Another snow storm is expected for tomorrow morning.

Note about the taumeter:

About focus offsets and the cold amplifier response of the Box C:

ar2_map_scan102

SB: Nice! It looks great! Thank you! Also, the difference in focus between Ar2 and Ar1 does not see to have changed since January, i.e. after the new full warm up. Also good news.

Friday, 13th February 2026

00:00-06:00: Maite. A new storm is coming. Snowfall/frozen rain episodes and very strong winds (estimated 30+ m/s). No observations.

06:00-13:00: Philippe. Adverse weather continues. No observations.

13:00-19:00: David. More adverse weather. No observations.

19:00-23:00: Angel. Light snowfall. Hail. Dense fog.

23:00-24:00: Maite. Frozen rain, strong winds. Deicing ON (02:00).

Saturday, 14th February 2026

00:00-06:00: Maite. Snow and freezing rain continue, along with high-speed winds.

06:00-13:00: Philippe. The adverse weather continues unchanged. Deicing is still active (high power; fortunately, the cost per kWh during nights and weekends corresponds to the off-peak tariff).

telescope view 20260214 morning control room access

forecast 20260214 10:45 UT

13:00-19:00: David. Wind speed has been dropping, dense fog all the time, but atmospheric pressure is steeply rising (by the way, the famous Azores High regains its seasonal geographic position, at last). Tighter monitoring of the glycolated mix temperature of the NIKA2 He compressors by the Operators.

19:00-23:00: Angel. Foggy, high clouds. High winds now from NW direction. Deicing is still active.

telescope view 20260214 evening

23:00-24:00: Maite.

Sunday, 15th February 2026

00:00-06:00: Maite. Windy all night. Deicing continues.

06:00-13:00: Philippe. Clear and pretty skies. Too much windy for any critical operation outside/moving the telescope: ~15-20 m/s (OSN's anemometer came back to life).

13:00-23:00: David and Angel. Started at 1300, DAQ started in total power, bad kids array 1,2,3 respectively 15, 4, 29. Checking pointing behaviour with several scans. Checking that the homology corrections are correct. FWHM of point-like sources with best focus in array 1 showed a distorted beam. Checked the temperature differences of the backup structure and quadrupod wrt. yoke-reference are far from being stable, so the first focus sequences are not confident (only the last two in this session were recorded in Table above).

23:00-24:00: Maite. Only 3 scans of the map of 147-25 done. Scan errors of 0.1% to 10.5%, due to tracking deviations greater than 3 arcsec (wind gusts). Telescope parked at 23:15 due to high winds.

Monday, 16th February 2026

00:00-06:00: Maite. The wind has dropped. tau_225GHz ~0.4. Starting again at 00:45 with EMIR project 102-24. Nearby calibrators are very weak. Despite using the faint calibrators script, need to do different pointings before obtaining a reasonable one. Problems with the tuning after 1.15h. Starting again from the beginning at 02:00. Tracking alarm set a couple of times despite no wind. 2 repetitions. At 3:30 the opacity is pretty low (~0.17), but the wind has increased and it is given some track errors.

06:00-13:00: Philippe. Low opacity (< 0.2), stable conditions (better than +/- 5 Jy/beam Ar2) at start of shift.

13:00-22:00: David and Angel. Moderate opacity (~0.4-0.5, above max. values allowed for NIKA2 programs), so stayed on EMIR.

22:00-24:00: Project 175-25 (EMIR). Observations OK. Excellent weather.

Tuesday, 17th February 2026

00:00-02:00: Project 175-25 (EMIR). Observations OK.

02:00-07:00: 087-21 (EMIR/XPOL). Observations OK.

07:00-10:30: Project 175-25 (EMIR). Observations OK.

10:30: Maintenance time.

The 5th week of the NIKA2 winter pool ends. Thanks to Gloria and Pedro for their invaluable support in operations, which extends to the entire IRAM technical staff, as well as to Aida, Sara, and Alba for helping us keep morale high under adverse weather with their fabulous meals, and to the observers Maite, Philippe, David, and Miren for their kind participation. Hope to see you again.

last shot

DailyReportsNika2Pool100226 (last edited 2026-02-18 11:39:04 by NikaBolometer)